if the point (-4,k) is on the graph of the equation 3x+y=-8 find the value of k
Found 2 solutions by richwmiller, unlockmath:
Answer by richwmiller(17219)

if x=-4 what is y?
3(-4)+y=-8
-12+y=-8
y=4
there fore k=4

Answer by unlockmath(1688)

Hello,
First, we can plug -4 for "x" into the equation.
3x+y=-8
3(-4)+y=-8 Then add 12 to both sides to give us:
y=12-8 now combine like terms:
y=4 Therefore:
k would be 4
Sometimes letters confuse students but k here is showing the y value.
RJ
Check out a new book I wrote at:
www.math-unlock.com